What is the main type of power plant using in Oman? What are the modifications that we can do to improve the performance of the plant? Explain with diagrams. How it will improve the performance?
Answer #1

Due to abundace of solar light in arab countries , the major source of power can be Sun there . Hence , majorly it uses:

1. Solar Power Plants ( SOHAR ALIMINIUM POWER PLANT with a capacity of 1000MWE )

2. Gas Power Plants ( Barka II CCGT Power and Desalination Plant with a capacity of 728 MWE )

MODIFICATIONS CAN BE DONE FOR IMPROVING PERFORMANCE :

1. For Solar Power Plants :

1. Using High efficiency PV cells .

2. By decreasing the cost per area of the cells.

3. Use Of Proper Solar tracking Systems .

4. Increasing the absorptive efficiency of the selective coating applied on the Cell.

5. Using Efficient Storage Devices .

2. For Gas Power Plants :

1. Regeneration in Gas turbines .

2. Reheating of the steam .

3. Intercooling in Gas Turbines .

Caselo REGENERATION in GAS TURBINE; It is the preolen in whin wroine exhaust all utilised to f the tempo of Compressed air be>wwwww KURBINE I Ty > Tz Combacho -SZO 11 I Tu> Ta> Tz 1-2-a- 3 - bry 1-2 -3-4 I We = 12 - hi 2 Wy z hz-hy I we=hach, Wed ny-- 3 CASE-2 REGGEATING The main purpose of Reboting is to by ping the tuebue worst Poster T> Та 2 т.1-2-34-5-6-1 (Rou) Wes ne che const. we is grha u fur =( hg - ha) # (ushe) Neste wawe of Whet af Wp - We Vo = ng-h2. 1 * cl=88 3: INTER COOLING IN G.TR ug another alay of improving the net wolk by dog the comprenor work. INTERCOOLER S ROOLER & Sce J